The spiral spring of a spring balance is 25cm, when 5 hang on it and 30cm long when the wait is 10N. What is the length of the spring if the weight is 3N
1
Expert's answer
2020-12-14T07:17:10-0500
"W_1=k(x_1-x_0)\\W_2=k(x_2-x_0)\\\frac{x_2-x_0}{x_1-x_0}=\frac{W_2}{W_1}
\\\frac{30-x_0}{25-x_0}=\frac{10}{5}\\x_0=20\ cm\\5=k(25-20)\\k=1\frac{N}{cm}"

Thus,

"3=1(x-20)\to x=23\ cm"